OpenStudy (anonymous):

Simplify: 4a + 3ab + 6ab

OpenStudy (unklerhaukus):

\[4a + 3ab + 6ab=4a+9ab=a(4+9b) \]

OpenStudy (anonymous):

always combine similar terms

OpenStudy (anonymous):

4a + 3ab + 6ab = add the like terms i.e 4a + 3ab +6ab => 4a + 9ab

OpenStudy (anonymous):

=a(4+3b+6b)

OpenStudy (anonymous):

@pastory you forgot to combine 3b and 6b

OpenStudy (anonymous):

=a(4+9b) yes yes TRALLE.